Circular No. 6339

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744–28 A. J. Castro-Tirado, Laboratorio de Astrofísica Espacial y Física Fun-

damental, Madrid; R. Gonzalez-Delgado, Space Telescope Science Institute; E. Perez, Instituto de Astrofísica de Andalucia, Granada; and J. M. Rodriguez-Espinosa, Instituto de Astrofísica de Canarias, report: “On Mar. 3 and 5, we observed the entire XTE error box of GRO J1744–28 (IAUC 6309) with the 3.5-m telescope (+ MAGIC in high-resolution mode, 0′′.31 per pixel) at the German-Spanish Calar Alto Observatory. Images were taken using two narrow-band filters centered at 2.166 μm (Brackett-γ, width 0.022 μm) and 2.26 μm (continuum K, width 0.060 μm). Co-added images (total exposures 1560 s at 2.166 μm and 250 s at 2.26 μm) were obtained in order to search for Brγ emission-line objects within the XTE error box. No such objects brighter than K = 14 were found within the XTE error box or at the position of the variable radio source (IAUC 6307, 6323) that has been recently confirmed by ASCA (IAUC 6337) as the radio counterpart of GRO J1744–28. Taking into account the optical/infrared astrometry (IAUC 6315, 6318) and the large N_{H} value derived by ASCA, it seems that ‘star C’ (K = 13.3, IAUC 6318, 6321) is a foreground star ∼ 0′′.5 away from GRO J1744–28 (similar to GX 17+2, whose optical counterpart is almost coincident with the position of a foreground G star; Naylor et al. 1991, MNRAS 252, 203). For a source distance of 7 kpc, we estimate 13.5 < K < 17.5. Therefore, only deep, high-resolution, and preferably thermal imaging of GRO J1744–28 while in outburst may reveal the infrared counterpart.”

SUPERNOVA 1993J IN NGC 3031 H.-U. Zimmermann and B. Aschenbach, Max-Planck-Institut für Ex-

traterrestrische Physik; and W. H. G. Lewin, Massachusetts Institute of Technology, report: “SN 1993J in M81 has been monitored twice a year with the HRI detector onboard the x-ray satellite ROSAT. The latest measurements from Oct. 1995 confirm that, after the strong initial decrease of the soft x-ray flux during the first year, the soft x-ray flux has been continuously rising since early 1994. Assuming a temperature of kT = 1.25 keV and foreground absorption of N_{H} = 5.4 × 10^{20} cm^{−2} as deduced from a Raymond-Smith fit to PSPC data taken at the end of 1993, the absorptioncorrected luminosity in the ROSAT band (0.1–2.4 keV) in Oct. 1995 is 5.4 × 10^{38} erg/s, which is ∼ 70 percent higher than in Apr. 1994.”

1996 March 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green